### v3

# Reverse Engineering the ESP32-C3 Wi-Fi Drivers for Static Worst-Case Analysis of Intermittently-Powered Systems

Abstract.
The Internet of Batteryless Things revolutionizes sustainable communication as it operates on harvested energy. This harvested energy is dependent on unpredictable environmental conditions; therefore, device operations, including those of its networking stack, must be resilient to power failures. Reactive intermittent computing provides an approach for solving this by notifications of impending power failures, which is implemented by monitoring the harvested energy buffered in a capacitor. However, to use this power-failure notification and guarantee forward progress, systems must break down tasks into atomic transactions that can be predictably finished before the energy runs out. Thus, static program-code analysis must determine the worst-case energy consumption (WCEC) of all transactions. In Wi-Fi–capable devices, drivers are often closed-source, which avoids the determination of WCEC bounds for transactions since static analysis requires all code along with its semantics.
In this work, we integrate an energy-aware networking stack with reverse-engineered Wi-Fi drivers to enable full-stack WCEC analysis for physical transmission and reception of packets. Further, we extended a static worst-case analysis tool with a resource-consumption model of our Wi-Fi driver. Our evaluation with the RISC-V–based ESP32-C3 platform gives worst-case bounds with our static analysis approach for the transactions of the full communication stack, therefore showing that Wi-Fi–based reactive intermittent computing is feasible.
1. Introduction
Batteryless Systems
Batteryless devices (Ahmed et al., 2024) rely on harvested energy. Being batteryless not only results in independence from battery requirements but also makes novel energy-self–sufficient applications possible: Batteryless medical implants (Kim et al., 2024) reduce the need for surgical replacements due to degradation in battery performance, putting them into the scope of safety- and security-critical systems (Mottola et al., 2024). Naturally, the unpredictable nature of harvesting energy means that a started task can face a power failure before reaching a checkpoint stored in non-volatile memory. Power failures cause two major problems: First, failures can leave the system in an inconsistent state. Second, with failures during energy-intensive operations with transactional semantics (e.g., sending packets), systems might repeatedly face failures without making progress.
Worst-Case Analysis
To address inconsistency and starvation, static worst-case program analysis can guarantee power-failure–free execution in intermittent systems. Specifically, static worst-case energy consumption (WCEC) analysis (Choi et al., 2022; Raffeck et al., 2024; Wägemann et al., 2018; Yarahmadi and Rohou, 2020) determines a safe upper bound on operations’ energy demand. These analysis techniques originally stem from determining the worst-case execution time (WCET) to guarantee timeliness in real-time systems (Wilhelm et al., 2008). With WCEC estimates and the available energy in the storage (i.e., capacitor) during runtime, operations execute without the risk of power failures. Consequently, consistency and forward progress are guaranteed under consideration of the currently available energy.
Available Code & Semantics
A necessary requirement for such static resource-consumption analyses is the availability of all code along with its resource-consumption semantics (Abella et al., 2015; Schoeberl, 2009). This knowledge is particularly important for power-consuming operations, such as radio transceivers, since inaccuracies in the static behavior modeling can result in unsound or impractically pessimistic analysis estimates. Unfortunately, for our target platform, the RISC-V–based ESP32-C3 system-on-chip (SoC), the Wi-Fi drivers are closed-source, and the Wi-Fi peripheral’s behavior is undocumented.
Contributions
The ESP32-Open-MAC project (Devreker, 2023) provides reverse-engineered Wi-Fi drivers for an Xtensa-based SoC, which we leveraged for the RISC-V–based ESP32-C3. Our work presents open-source Wi-Fi drivers for the ESP32-C3 alongside their static worst-case analyzability. In summary, our contributions are threefold:
Case Study on Reverse Engineering: We present our reverse engineering efforts for the Wi-Fi driver of the ESP32-C3.
Energy-Consumption & Execution-Time Model for Worst-Case Analysis: Our static analysis comes with a fine-grained energy-consumption model.
Open-Source Implementation & Evaluation: Using the reverse-engineered knowledge, we develop an open-source Wi-Fi driver for the ESP32-C3 and present evaluation results that validate the driver’s property of being statically analyzable.
2. Background
This section gives background on network stacks and mechanisms for processor and Wi-Fi module communication. Finally, we outline worst-case resource-consumption analyses of program code.
2.1. Network Stack
The network stack on embedded devices typically consists of a lightweight TCP/IP implementation and a driver that interacts with the physical network interface (e.g., Wi-Fi peripheral). At the highest level of the stack is the application that provides data to be transmitted, while at the lowest level is the physical layer (PHY), which sends out the data onto the physical medium. On top of the physical layer is the medium access control (MAC) that coordinates transmission between multiple devices. In the case of Wi-Fi, the carrier sense multiple access/collision avoidance (CSMA/CA) mechanism is used to coordinate transmissions. Each layer of the network stack encapsulates data from the upper layer and passes it to the layer below. On intermittent devices, power may be lost at any point before packet transmission is completed, leading to a loss of progress up to that point.
In the case of Wi-Fi, PHY-layer operations are implemented in hardware due to strict timing requirements. Some Wi-Fi peripherals implement MAC operations in software (SoftMAC), while others implement both MAC and PHY operations in hardware (FullMAC).
2.2. Wi-Fi Hardware
High-speed communication between the Wi-Fi peripheral and the processor within an SoC is typically carried out using a combination of memory-mapped I/O (MMIO) and direct memory access (DMA) (Schulz, 2018). The Wi-Fi peripheral is configured through MMIO, while data transfer for transmission and reception is performed using DMA. DMA allows the CPU to remain free for other tasks instead of handling the transfer. Espressif devices such as ESP32-C3 follow this architecture. The peripheral communicates events such as the successful transmission or reception of a packet through interrupts. On other platforms, external Wi-Fi chipsets may use the SPI interface to communicate with the host controller.
2.3. Static Worst-Case Resource Analysis
Static worst-case resource-consumption analysis of program code comprises two major pillars: (1) a hardware-agnostic program path analysis and (2) a hardware-dependent cost modeling capturing the actual resource demand (i.e., execution time or energy demand) for the actual target platform. The main idea is to formulate a mathematical problem formulation, usually an integer linear program (ILP): The ILP’s constraints originate from the path analysis. Abstract-interpretation techniques are further useful for refining the path constraints in a context-sensitive manner. The ILP’s objective function states the maximization of the actual costs with the given path constraints. Eventually, the ILP’s solution yields a mathematically proven worst-case bound. With regard to energy consumption, the WCEC bound is useful to guarantee the execution without power failures during the runtime of batteryless devices.
3. Problem Statement
The problems that we address in this work are twofold: Section 3.1 discusses the problem of closed-source drivers, and Section 3.2 explains the requirement of resource models.
3.1. Problem of Closed-Source Wi-Fi Drivers
In intermittently-powered devices, static worst-case analysis techniques can be crucial to gain knowledge on the worst-case energy consumption of operations (Choi et al., 2022; Raffeck et al., 2024; Yarahmadi and Rohou, 2020). During runtime, the system’s scheduler checks the available energy (comparable to just-in-time checkpointing) and reactively dispatches tasks only if sufficient energy is available to guarantee freedom from power failures. The WCEC-aware approaches give both memory-consistency and forward-progress guarantees. The trade-off we exploit favors expensive design-time analyses over complicated runtime checkpointing mechanisms. Once deployed, systems achieve provable resilience against power failures. However, these design-time analyses require the code to be available. Our target platform, the ESP32-C3, is a RISC-V–based, widely used SoC from Espressif, which has closed-source Wi-Fi drivers. Although less power-intensive networking technology exists (e.g., BLE), we are interested in the Wi-Fi stack being of high relevance: In 2024, an estimated 31 % of IoT devices used Wi-Fi for connectivity (Sinha, 2024). Further, the recent article on The Internet of Batteryless Things (Ahmed et al., 2024) calls for action and states that popular network protocols, including Wi-Fi, should be adapted for intermittent operation. Unfortunately, for our target, the generic Wi-Fi drivers were not available as analyzable source code, which is a common problem in embedded SoCs. Of main interest for our analysis are the MAC and PHY layers.
Our Approach in a Nutshell: To solve the problem of the closed source, we exploit several reverse-engineering techniques on Espressif’s closed-source Wi-Fi driver alongside utilizing existing knowledge (Devreker, 2023). Fortunately, Espressif legally allows reverse engineering this driver. Based on the reverse-engineering results, we contribute an open-source driver for the ESP32-C3.
3.2. Missing Resource-Consumption Model
Although the developed open-source driver gives us the freedom to perform static analysis, a major problem is connecting memory-mapped IO operations to semantics with regard to the power demand.
When considering the following MMIO operation:
*((uint32_t*)0x60033ca0) &= 0xff00efff,
the static analysis is not aware of its semantics. On the ESP32-C3, this operation switches the transceiver from a sleep state to standby mode (see also Figure 2).
Hence, our static resource analysis needs to include the notion that the power drawn increases after this operation.
Similarly, specific instructions also require a longer execution time than a normal RAM access, which in turn leads to higher energy demand (i.e., power over time).
Consequently, a comprehensive model needs to account for power, energy, and time.
Our Approach in a Nutshell: For our target, we present a resource-consumption model that considers transitions in the hardware, which change the SoC’s power demand. Our analysis makes use of this model and tracks program paths in a context-sensitive way.
4. Approach
We use a combination of commonly available static and dynamic reverse-engineering analysis methods for identifying the registers and, correspondingly, their contents.
4.1. Reverse Engineering
Static Analysis
Through decompilation, the program flow and, accordingly, the registers used for transmission and reception can be identified. Functions within the binaries are named based on their purpose, which leads to informative decompilation results. Additionally, the undocumented interrupt configuration used by the Wi-Fi peripheral can be identified. However, decompilation alone is insufficient to identify control and data flows. Several of the registers are configured based on data that is internal to the driver, whose structure is not documented. Moreover, the peripheral also has several slots for transmission, which are dynamically selected based on the network configuration. Therefore, we use QEMU to dynamically determine the register contents as well as the slots.
Dynamic Analysis
The QEMU version in (lcgamboa, 2023), a fork of Espressif’s QEMU, supports Wi-Fi for ESP32-C3 by emulating an open Wi-Fi access point (AP). In particular, it only implements registers required for triggering a transmission and reading received packets through DMA as well as the status register, which encodes the reasons for an interrupt occurrence. We further modify this fork to log all register accesses and their contents within the Wi-Fi peripheral’s address range. While the QEMU fork provides insights into accesses of the peripheral’s registers and their contents used during the emulation, it does not fully emulate the hardware. Hardware initialization, for example, is not emulated. Moreover, it is not always possible to deduce the purpose of the register operation.
Binary Hooking for Injecting Tracing Code
The modules of Espressif’s Wi-Fi driver are provided as archives (.a) of object (.o) files. These object files have undefined symbols referring to functions defined in other modules, such as the binary blobs of the ESP32-C3 ROMs. Their definition is resolved during the final linking of the executable. To trace the execution, we hook into these symbols by replacing the symbol names and exploiting the original symbol names in our custom firmware to identify the control flow.
These reverse-engineering methods, in combination with the existing infrastructure in ESP32-Open-MAC, yield the necessary information for constructing the Wi-Fi driver for the ESP32-C3.
4.2. The Open-Source ESP32-C3 Driver
Once the registers and transmit/receive mechanisms are identified, we use the architecture shown in Figure 1 to develop the driver.
MAC Task
The MAC task accepts packets in Ethernet (802.3) format from the IP layer. It adds the Wi-Fi MAC header and passes it to the driver for transmission. On the other hand, packets arriving from the driver are passed to the IP layer through a callback (CB). The MAC task provides a transmit primitive for the IP stack to queue packets for transmission. Conversely, a receive callback has to be provided by the IP stack.
Driver Task
The driver task interacts with the MMIO registers and manages memory allocated to DMA structures required by the peripheral. All memory is statically allocated to aid WCET/WCEC analysis. Processing of interrupts is deferred by the interrupt handler to the driver task. Incoming 802.11 packets are passed to the MAC task for further processing through a queue. The interrupt handler notifies the driver of the occurrence and reason for the interrupt using a hardware event queue. The MAC task also writes packets to be transmitted into this queue. On the other hand, the driver task writes any received packets into the MAC event queue to signal the MAC task of received packets.
[architecture]architecture of firmware
4.3. Building a Resource-Consumption Model
For device operations such as transmission and reception, the Wi-Fi peripheral operates independently of the CPU. To yield sound results, static resource analysis requires a cost model specifying the maximum time and energy demands of operations. These paragraphs provide a generic overview of the cost model, while Section 5.2 discusses concrete values.
Timing Information
The necessary timing-related costs comprise both protocol- and hardware-specific aspects. The application-specific protocol influences the maximum duration of device operations (e.g., through the maximum packet size). Additionally, timeouts and operation windows in the protocol allow deriving an upper bound on operations’ execution times, such as waiting to receive a packet. Bounds on the duration of operations also enable bounding the occurrence of interrupts related to the completion of a transmission or an incoming packet. The second aspect, the hardware-specific behavior, concerns the Wi-Fi peripheral’s MMIO interface. The execution time of certain MMIO operations, such as waiting for a bit to change its value, which signals the completion of a configuration update, depends on the reaction of the hardware.
[device graph]device graph showing energy states
Energy Information
The energy-related cost model for the Wi-Fi peripheral consists of all possible device states with distinct power consumption. A simplified version of the device graph modeling the states and their transitions for the Wi-Fi peripheral is shown in Figure 2. The depicted graph features three states of operation: Sleep, Standby, and Transmitting. The Wi-Fi peripheral is powered down and inoperative in the sleep state ( ). In the standby state ( ), the Wi-Fi peripheral is powered up and may receive packets. Lastly, in the transmitting state ( ), the device actively transmits a packet. No reception occurs in this state, as the peripheral owns the channel through the CSMA/CA mechanism. For this paper, we assume that the CPU is always on and uses only one run mode, although modeling the ESP32-C3’s sleep states, run modes, and clock-frequency changes is generally possible (Dengler et al., 2023).
4.4. Implementation
Reverse Engineering
The ESP32-Open-MAC project lists the necessary MMIO registers to achieve basic transmission and reception on the ESP32. On the ESP32, the MMIO addresses belong to the 0x3ff73000-0x3ff75fff range, but on the ESP32-C3 are mapped to 0x60033000-0x60035fff. The offsets of ESP32-C3’s MMIO registers from the base address differ from those of the ESP32. In order to identify the equivalent registers on the ESP32-C3, we write a firmware that performs simple UDP transmission and analyze it using Ghidra’s (Agency, 2019) decompiler. In Ghidra, we mark the memory ranges of the peripheral identified with QEMU (lcgamboa, 2023) as volatile and identify the functions that access these memory locations. Functions with prefixes such as hal_mac_* or wdev_mac_* were found to handle MAC/PHY layer operations, which are of particular interest for our reverse engineering.
Among the identified functions, we hook into those that are listed as undefined symbols in object files within the libpp.a archive of the proprietary blobs. The resulting firmware with the hooks is loaded into our QEMU fork.
| Register | Deduced Purpose | Access |
| 0x60033000 | MAC address configuration | R/W |
| 0x60033084 | MS bit enables or disables Rx | R/W |
| 0x60033088 | Set base DMA linked list address | R/W |
| 0x60033ca0 | Power up and power down Wi-Fi module | R/W |
| 0x60033d08 | Configure Tx DMA address and trigger Tx | R/W |
| 0x60033c3c | Get interrupt reason | R |
| 0x60033c40 | Clear interrupt | R/W |
| 0x60033cac | Clear Tx slot | R/W |
An excerpt of the most important registers identified through our reverse engineering is shown in Table 1, along with the deduced purpose of these registers.
Interrupt Handling
The details of the Wi-Fi interrupt handling are not exposed to developers. Through decompilation, we find that the Wi-Fi driver uses CPU interrupt number 1, which is marked as reserved in the manual (Espressif Systems, 2025b), and uses an undocumented register for that interrupt source at address 0x600c2000. With this information, registering a Wi-Fi interrupt handler follows the usual procedure on the ESP32-C3. First, the (undocumented) interrupt source register (0x600c2000) is cleared, and the interrupt is disabled by clearing bit 1 of the (documented) interrupt-enable register (0x600c2104). The interrupt handler is registered using ESP-IDF’s intr_handler_set. Finally, the interrupt is enabled by setting bit 1 of 0x600c2104 and loading the interrupt number to 0x600c2000.
ESP32-C3 Wi-Fi Driver
We implement the driver using the architecture shown in Figure 1. We use an open AP, as the encryption functionality of the Wi-Fi hardware is not fully reverse-engineered. For initialization and connecting to an AP, we use ESP-IDF’s API. The functions esp_wifi_init and esp_wifi_start initialize and configure the Wi-Fi, set the internal interrupt handler for Wi-Fi, and initiate an AP connection. Their successful completion can be monitored through the system notification WIFI_EVENT_STA_CONNECTED. Once the notification is received, the Wi-Fi task can be killed using the undocumented pp_post function, which queues instruction codes to the Wi-Fi task started by the driver. Finally, our driver and MAC tasks are started. We use PfIP (Vogelgesang et al., 2025) as the IP stack due to its intermittency-aware execution.
5. Evaluation
We perform a static resource analysis of our Wi-Fi driver to assess the ability to bound the program flow and resource demand. Section 5.1 describes the hardware and network setup as well as the task system used for the experiments. Additionally, we detail the required assumptions for the static resource-consumption analysis using the Platin analysis framework (Maroun et al., 2024). In Section 5.2, we examine and discuss the analysis results.
5.1. Evaluation Setup
Hardware Setup
[…]
Functionality
To validate the driver’s functionality, we write a firmware that sends UDP packets through PfIP’s network stack. PfIP is provided with the mac_task’s transmit handle, which in turn queues the packet for physical transmission by the hw_task. The responses from the network are passed back through PfIP’s receive callback. We capture the network traffic using Wireshark to monitor our driver’s functional correctness.
Static Resource Analysis
To analyze the worst-case resource demand of our Wi-Fi driver, we use the Platin analysis framework (Maroun et al., 2024). With the device-model analysis of the WoCA approach (Raffeck et al., 2024), Platin supports device-aware resource-consumption analysis for the ESP32-C3. Platin uses a modified version of the clang compiler that collects user-annotated knowledge, derives metainformation about the program flow, and tracks the relation between source and object level regarding this control-flow information. Using this metainformation and hardware-specific cost models, Platin formulates the costs of and constraints on the program behavior as an ILP suitable for common mathematical solvers, such as lp_solve (Berkelaar et al., 2004) or Gurobi (Gurobi Optimization, LLC, 2025), to obtain the maximum resource demand. For the purpose of this paper, we annotate the current consumption for each state in the device’s energy model based on the values provided by the manufacturer in the datasheet (Espressif Systems, 2022). For the processor, we rely on the model for the ESP32-C3 already present in Platin.
The focus of our evaluation lies on the driver functions, independent of the runtime and operating system. While a whole-system analysis, which includes the operating system, is generally possible (Schuster et al., 2019), we exclude the FreeRTOS components and callbacks into the PfIP network stack from our analysis for this paper.
We analyze benchmarks including individual driver functions and an evaluation task comprising multiple device interactions. For each driver function, we determine an upper bound on the execution time. The evaluation task sends a packet and actively waits for the transmission’s completion. After starting the transmission, the Wi-Fi transceiver transmits the packet and waits for an acknowledgment (ACK). Once the transmission is completed, the transmission returns to standby mode. We derive an upper bound on both the time and energy demand with respect to the different device states of the Wi-Fi transceiver during the task execution.
[…]
5.2. Evaluation Results
| WCET | always-on WCEC | ||
| wifi_hw_deinit | 48 cy | 0.3 us | 0.3099 uJ |
| wifi_setup_interrupt | 178 cy | 1.11 us | 1.149 uJ |
| wifi_setup_rx | 1881 cy | 11.8 us | 12.14 uJ |
| wifi_hw_init | 49 cy | 0.306 us | 0.3163 uJ |
| wifi_transmit_packet | 335 cy | 2.09 us | 2.163 uJ |
| wifi_wait_for_tx | 52184 cy | 326 us | 336.9 uJ |
| wifi_process_tx_done | 157 cy | 0.981 us | 1.014 uJ |
| wifi_handle_rx | 12989 cy | 81.2 us | 83.85 uJ |
| wifi_process_timeout | 138 cy | 0.862 us | 0.8909 uJ |
| wifi_get_bssid | 94 cy | 0.588 us | 0.6068 uJ |
| wifi_mac_handle_rx | 68715 cy | 429 us | 443.6 uJ |
| wifi_interrupt_handler | 943 cy | 5.89 us | 6.088 uJ |
| WCET | always-on WCEC | device-aware WCEC | ||
| TX task | 52615 cy | 329 us | 339.7 uJ | 32.42 uJ |
[…]
Most functions of our open-source Wi-Fi driver perform only operations on memory-mapped registers to control the transceiver and exhibit a relatively low execution time and energy consumption. The few notably higher resource bounds relate to worst-case analysis assumptions for the control flow. wifi_mac_handle_rx moves the payload of the received package between buffers, so the analysis has to assume the worst case of the maximum packet size. Similarly, the analysis has to assume the worst-case number of packets for the execution time of wifi_handle_rx and the maximum time until an ACK is received for wifi_wait_for_tx.
The difference between the always-on and device-aware WCEC bound underlines the importance of considering the states of the Wi-Fi transceiver in a fine-granular manner. We demonstrated that the reverse engineering of our driver code enables deriving the necessary control-flow bounds and performing accurate static resource analysis of Wi-Fi operations on the ESP32-C3.
6. Related Work
Intermittency-Aware Wireless Communication
Wireless networking is a key challenge in batteryless designs due to intermittency. Networking stacks deployed on such devices must be fully aware of intermittent execution (Fu et al., 2023). In contrast to our work on Wi-Fi, other approaches for batteryless wireless communication use Bluetooth Low Energy (BLE), LoRa, or backscatter-based wireless communications (Cai et al., 2023; Talla et al., 2017; Jiang et al., 2023). As a time-predictable IP stack, tpIP (Schoeberl and Pedersen, 2018) is designed to allow static analysis through bounded loops and avoids dynamic memory allocation. Similarly, Vogelgesang et al. designed PfIP, a networking stack with transactional semantics and worst-case bounds for every transaction (Vogelgesang et al., 2025). Transactional semantics break down the overall task into atomic sub-tasks (or transactions) that can be independent of each other. These approaches are agnostic to the underlying physical layer.
FreeBie (de Winkel et al., 2022) is a bidirectional BLE stack designed for intermittent communication, demonstrated with a batteryless smartwatch. Cameroptera (Nardello et al., 2019) is a LoRa-based batteryless image sensing system that adaptively configures its transmission mode based on the available energy. RockClimb (Choi et al., 2022) uses WCET bounds without device awareness to guarantee the reaching of checkpoints. Yarahmadi et al. (Yarahmadi and Rohou, 2020) employ WCEC-aware checkpoint placement without considering the resource demand of peripheral devices. WoCA introduces device-aware whole-system worst-case analysis for intermittent systems, including a model for LoRa-based communication (Raffeck et al., 2024).
Without an intermittency-aware protocol, two devices may not be awake at the same time for communication. Greentooth (Babatunde et al., 2024) synchronizes communication between two devices using time division multiple access (TDMA)-style scheduling and low-power wake-up radios. Probabilistic modeling of latency or charge times (Geissdoerfer and Zimmerling, 2021, 2022) likewise achieves synchronized communication.
For intermittent devices, Wi-Fi is under-explored due to its high power demand but promises batteryless innovation, for example, in fields like smart-home applications.
Open-Source Wi-Fi Platforms
Outside the world of low-power SoCs, open-source drivers exist for several Wi-Fi network interface cards, often supported by the manufacturers themselves, for example, from Atheros or Mediatek (Project, 2024). The OpenWiFi project (Jiao et al., 2020) developed an open-source FPGA-based software-defined radio for the 802.11 standard paired with an RF frontend.
On low-power devices, the ZeroWi project (Bentham, 2020) reverse-engineers the capabilities of the Cypress CYW43438 Wi-Fi SoC, which is programmed using non-free firmware. Drivers written for this chip then send commands in the SDIO format over SPI to communicate with the firmware. Early attempts at reverse engineering Espressif’s Wi-Fi driver were performed by Uri Shaked as part of the Wokwi simulator for ESP32 (Shaked, 2019). The QEMU-based open-source PICSimLab simulator introduces support for Wi-Fi emulation on the ESP32 and ESP32-C3 (lcgamboa, 2022). Based on these attempts, the ESP32-Open-MAC project (Devreker, 2023) aims to fully open-source the Wi-Fi driver on the ESP32. Our work not only ported ESP32-Open-MAC onto the RISC-V–based ESP32-C3 but also integrated it with a resource-consumption model for static analysis and a transactional network stack that is suitable for batteryless devices.
7. Future Work & Conclusion
The use of Wi-Fi for intermittently-powered batteryless systems is an ongoing research challenge since network protocols are not designed with intermittent operation in mind (Ahmed et al., 2024). With this work, we contribute to this challenge with our reverse-engineered and publicly-available Wi-Fi drivers, along with their static worst-case energy-consumption analysis. Our analysis results allow intermittent systems to execute with runtime guarantees on the power-failure–free execution of operations.
Future work needs to consider a more detailed device model to reduce the pessimism in the bounds. For example, in , the peripheral is not actively transmitting after the end of transmission until the ACK is received. Additionally, energy penalties for switching on/off energy states could be accurately considered. We also use a static peripheral configuration. The resource-consumption bounds are impacted by peripheral configurations, such as data rate of transmission. Future work could identify optimal configurations under energy constraints. ESP32-Open-MAC has recently introduced an initialization of the Wi-Fi peripheral independent of FreeRTOS. Therefore, an OS-aware implementation could also result in easier static analysis, for example, by exploiting Rust’s statically analyzable async/await pattern.
In this work, we presented a fully open-source Wi-Fi driver for the ESP32-C3, based on the ESP32-Open-MAC project. We successfully port the project from the Xtensa-based ESP32 to the RISC-V–based ESP32-C3 using common reverse-engineering techniques. Finally, we present a proof-of-concept for performing static resource-consumption analysis on the Wi-Fi peripheral’s physical operations using the extended Platin worst-case analysis toolkit.
[ backgroundcolor=tolerableGreen!25, linecolor=tolerableGreen,linewidth=2pt,topline=false,bottomline=false, innertopmargin=0pt, innerbottommargin=3pt, innerleftmargin=6pt, innerrightmargin=6pt, frametitleaboveskip=3pt, frametitlefont=, frametitlealignment=, frametitle=The implementation is available at: ] https://git.nt.uni-saarland.de/open-access/pfip
Acknowledgements.
This work is supported by the German Research Foundation (DFG) as part of SPP 2378 (Resilient Worlds: project number 502615015, Resilient Power-Constrained Embedded Communication Terminals, ResPECT) and the NGI0 Core Fund under grant agreement No 101092990. We would also like to thank all ESP32-Open-MAC contributors for their valuable contributions to the community.References
- (1)
- Abella et al. (2015) Jaume Abella, Carles Hernández, Eduardo Quiñones, Francisco J. Cazorla, Philippa Ryan Conmy, Mikel Azkarate-askasua, Jon Pérez, Enrico Mezzetti, and Tullio Vardanega. 2015. WCET analysis methods: Pitfalls and challenges on their trustworthiness. In Proceedings of the 10th International Symposium on Industrial Embedded Systems (SIES ’15). 39–48. https://doi.org/10.1109/SIES.2015.7185039
- Agency (2019) National Security Agency. 2019. Ghidra. https://ghidra-sre.org/.
- Ahmed et al. (2024) Saad Ahmed, Bashima Islam, Kasim Sinan Yildirim, Marco Zimmerling, Przemyslaw Pawelczak, Muhammad Hamad Alizai, Brandon Lucia, Luca Mottola, Jacob Sorber, and Josiah D. Hester. 2024. The Internet of Batteryless Things. Commun. ACM 67, 3 (2024), 64–73. https://doi.org/10.1145/3624718
- Babatunde et al. (2024) Simeon Babatunde, Arwa Alsubhi, Josiah D. Hester, and Jacob Sorber. 2024. Greentooth: Robust and Energy Efficient Wireless Networking for Batteryless Devices. ACM Transactions on Sensor Networks 20, 3 (2024), 66:1–66:31. https://doi.org/10.1145/3649221
- Bentham (2020) Jeremy Bentham. 2020. Zerowi: Bare-Metal WiFi Driver for the Raspberry Pi. https://iosoft.blog/zerowi/
- Berkelaar et al. (2004) Michel Berkelaar, Kjell Eikland, and Peter Notebaert. 2004. lp_solve 5.5, open source (mixed-integer) linear programming system. Software.
- Cai et al. (2023) Zhipeng Cai, Quan Chen, Tuo Shi, Tongxin Zhu, Kunyi Chen, and Yingshu Li. 2023. Battery-Free Wireless Sensor Networks: A Comprehensive Survey. IEEE Internet of Things Journal 10 (2023), 5543–5570. https://doi.org/10.1109/JIOT.2022.3222386
- Choi et al. (2022) Jongouk Choi, Larry Kittinger, Qingrui Liu, and Changhee Jung. 2022. Compiler-Directed High-Performance Intermittent Computation with Power Failure Immunity. In Proceedings of the 28th Real-Time and Embedded Technology and Applications Symposium (RTAS ’22). 40–54. https://doi.org/10.1109/RTAS54340.2022.00012
- de Winkel et al. (2022) Jasper de Winkel, Haozhe Tang, and Przemyslaw Pawelczak. 2022. Intermittently-Powered Bluetooth That Works. In Proceedings of the 20th Annual International Conference on Mobile Systems, Applications and Services (MobiSys ’22) (Portland, Oregon) (MobiSys ’22). 287–301. https://doi.org/10.1145/3498361.3538934
- Dengler et al. (2023) Eva Dengler, Phillip Raffeck, Simon Schuster, and Peter Wägemann. 2023. FusionClock: Energy-Optimal Clock-Tree Reconfigurations for Energy-Constrained Real-Time Systems. In Proceedings of the 35th Euromicro Conference on Real-Time Systems (ECRTS ’23), Vol. 262. 6:1–6:24. https://doi.org/10.4230/DARTS.9.1.2
- Devreker (2023) Jasper Devreker. 2023. Unveiling secrets of the ESP32: creating an open-source MAC Layer. Retrieved 2025-01-03 from https://zeus.ugent.be/blog/23-24/open-source-esp32-wifi-mac/
- Espressif Systems (2022) Espressif Systems. 2022. ESP32-C3 Series Datasheet. https://www.espressif.com/sites/default/files/documentation/esp32-c3_datasheet_en.pdf
- Espressif Systems (2025a) Espressif Systems. 2025a. ESP IoT Development Framework. https://www.espressif.com/en/products/sdks/esp-idf
- Espressif Systems (2025b) Espressif Systems. 2025b. ESP32-C3 Technical Reference Manual. https://www.espressif.com/sites/default/files/documentation/esp32-c3_technical_reference_manual_en.pdf
- Fu et al. (2023) Shen Fu, Vishak Narayanan, Mathew L. Wymore, Vishal Deep, Henry Duwe, and Daji Qiao. 2023. No battery, no problem: Challenges and opportunities in batteryless intermittent networks. Journal of Communications and Networks 25, 6 (2023), 806–813. https://doi.org/10.23919/jcn.2023.000033
- Geissdoerfer and Zimmerling (2021) Kai Geissdoerfer and Marco Zimmerling. 2021. Bootstrapping Battery-free Wireless Networks: Efficient Neighbor Discovery and Synchronization in the Face of Intermittency. In 18th USENIX Symposium on Networked Systems Design and Implementation (NSDI 21). 439–455. https://www.usenix.org/conference/nsdi21/presentation/geissdoerfer
- Geissdoerfer and Zimmerling (2022) Kai Geissdoerfer and Marco Zimmerling. 2022. Learning to Communicate Effectively Between Battery-free Devices. In 19th USENIX Symposium on Networked Systems Design and Implementation (NSDI 22). 419–435. https://www.usenix.org/conference/nsdi22/presentation/geissdoerfer
- Gurobi Optimization, LLC (2025) Gurobi Optimization, LLC. 2025. Reference Manual. gurobi.com
- Jiang et al. (2023) Tao Jiang, Yu Zhang, Wenyuan Ma, Miaoran Peng, Yuxiang Peng, Mingjie Feng, and Guanghua Liu. 2023. Backscatter Communication Meets Practical Battery-Free Internet of Things: A Survey and Outlook. IEEE Communications Surveys & Tutorials 25, 3 (2023), 2021–2051. https://doi.org/10.1109/COMST.2023.3278239
- Jiao et al. (2020) Xianjun Jiao, Wei Liu, Michael T. Mehari, Muhammad Aslam, and Ingrid Moerman. 2020. openwifi: a free and open-source IEEE802.11 SDR implementation on SoC. In Proceedings of the 91st IEEE Vehicular Technology Conference (VTC 2020). IEEE, 1–2. https://doi.org/10.1109/VTC2020-Spring48590.2020.9128614
- Kim et al. (2024) Young-Jun Kim, Jiho Lee, Joon-Ha Hwang, Youngwook Chung, Byung-Joon Park, Junho Kim, So-Hee Kim, Junseung Mun, Hong-Joon Yoon, Sung-Min Park, et al. 2024. High-Performing and Capacitive-Matched Triboelectric Implants Driven by Ultrasound. Advanced Materials 36, 2 (2024), 2307194. https://doi.org/10.1002/adma.202307194
- lcgamboa (2022) lcgamboa. 2022. Picsimlab. https://lcgamboa.github.io/
- lcgamboa (2023) lcgamboa. 2023. Qemu: Modified Version of the Espressif QEMU Used by the PICSimLab Simulator. https://github.com/lcgamboa/qemu/tree/picsimlab-esp32
- Maroun et al. (2024) Emad Jacob Maroun, Eva Dengler, Christian Dietrich, Stefan Hepp, Henriette Herzog, Benedikt Huber, Jens Knoop, Daniel Wiltsche-Prokesch, Peter P. Puschner, Phillip Raffeck, Martin Schoeberl, Simon Schuster, and Peter Wägemann. 2024. The Platin Multi-Target Worst-Case Analysis Tool. In Proceedings of the 22nd International Workshop on Worst-Case Execution Time Analysis (WCET ’24), Vol. 121. 2:1–2:14. https://doi.org/10.4230/OASIcs.WCET.2024.2
- Mottola et al. (2024) Luca Mottola, Arslan Hameed, and Thiemo Voigt. 2024. Energy Attacks in the Battery-less Internet of Things: Directions for the Future. In Proceedings of the 17th European Workshop on Systems Security (EuroSec ’24). 29–36. https://doi.org/10.1145/3642974.3652283
- Nardello et al. (2019) Matteo Nardello, Harsh Desai, Davide Brunelli, and Brandon Lucia. 2019. Camaroptera: A batteryless long-range remote visual sensing system. In Proceedings of the 7th International Workshop on Energy Harvesting & Energy-Neutral Sensing Systems (EnsSys ’19). 8–14. https://doi.org/10.1145/3362053.3363491
- Project (2024) Linux Wireless Documentation Project. 2024. Existing Linux Wireless Drivers — Linux Wireless Documentation. https://wireless.docs.kernel.org/en/latest/en/users/drivers.html
- Raffeck et al. (2024) Phillip Raffeck, Johannes Maier, and Peter Wägemann. 2024. WoCA: Avoiding Intermittent Execution in Embedded Systems by Worst-Case Analyses with Device States. In Proceedings of the 25th ACM SIGPLAN/SIGBED International Conference on Languages, Compilers, and Tools for Embedded Systems (LCTES ’24). 83–94. https://doi.org/10.1145/3652032.3657569
- Schoeberl (2009) Martin Schoeberl. 2009. Time-predictable computer architecture. EURASIP Journal on Embedded Systems 2009 (2009), 1–17. https://doi.org/10.1155/2009/758480
- Schoeberl and Pedersen (2018) Martin Schoeberl and Rasmus Ulslev Pedersen. 2018. tpip: A time-predictable tcp/ip stack for cyber-physical systems. In 2018 IEEE 21st International Symposium on Real-Time Distributed Computing (ISORC). IEEE, 75–82. https://doi.org/10.1109/ISORC.2018.00018
- Schulz (2018) Matthias Schulz. 2018. Teaching Your Wireless Card New Tricks: Smartphone Performance and Security Enhancements Through Wi-Fi Firmware Modifications. Ph. D. Dissertation. Darmstadt University of Technology, Germany. http://tuprints.ulb.tu-darmstadt.de/7243/
- Schuster et al. (2019) Simon Schuster, Peter Wägemann, Peter Ulbrich, and Wolfgang Schröder-Preikschat. 2019. Proving Real-Time Capability of Generic Operating Systems by System-Aware Timing Analysis. In Proceedings of the 25th Real-Time and Embedded Technology and Applications Symposium (RTAS ’19). 318–330. https://doi.org/10.1109/RTAS.2019.00034
- Shaked (2019) Uri Shaked. 2019. Wokwi - World’s Most Advanced ESP32 Simulator. https://wokwi.com/
- Sinha (2024) Satyajit Sinha. 2024. State of IoT 2024: Number of Connected IoT Devices Growing 13% to 18.8 Billion Globally. https://iot-analytics.com/number-connected-iot-devices/.
- Talla et al. (2017) Vamsi Talla, Bryce Kellogg, Shyamnath Gollakota, and Joshua R. Smith. 2017. Battery-Free Cellphone. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies 1, 2 (2017), 25:1–25:20. https://doi.org/10.1145/3090090
- Vogelgesang et al. (2025) Kai Vogelgesang, Ishwar Mudraje, Luis Gerhorst, Phillip Raffeck, Peter Wägemann, Thorsten Herfet, and Wolfgang Schröder-Preikschat. 2025. PfIP: A UDP/IP Transactional Network Stack for Power-Failure Resilience in Embedded Systems. In Proceedings of the Consumer Communications & Networking Conference (CCNC 2025). https://www.nt.uni-saarland.de/publications/vogelgesang2025pfip/PfIP_authors_version.pdf
- Wägemann et al. (2018) Peter Wägemann, Christian Dietrich, Tobias Distler, Peter Ulbrich, and Wolfgang Schröder-Preikschat. 2018. Whole-System Worst-Case Energy-Consumption Analysis for Energy-Constrained Real-Time Systems. In Proceedings of the 30th Euromicro Conference on Real-Time Systems (ECRTS ’18), Vol. 106. 24:1–24:25. https://doi.org/10.4230/LIPIcs.ECRTS.2018.24
- Wilhelm et al. (2008) Reinhard Wilhelm, Jakob Engblom, Andreas Ermedahl, Niklas Holsti, Stephan Thesing, David B. Whalley, Guillem Bernat, Christian Ferdinand, Reinhold Heckmann, Tulika Mitra, Frank Mueller, Isabelle Puaut, Peter P. Puschner, Jan Staschulat, and Per Stenström. 2008. The Worst-case Execution-time Problem – Overview of Methods and Survey of Tools. ACM Transactions on Embedded Computing Systems (ACM TECS) 7, 3 (2008), 36:1–36:53. https://doi.org/10.1145/1347375.1347389
- Wireshark Foundation (2025) Wireshark Foundation. 2025. free and open-source packet analyzer. https://www.wireshark.org/.
- Yarahmadi and Rohou (2020) Bahram Yarahmadi and Erven Rohou. 2020. Compiler optimizations for safe insertion of checkpoints in intermittently powered systems. In Proceedings of the 20th International Conference on Embedded Computer Systems: Architectures, Modeling, and Simulation (SAMOS ’20), Vol. 12471. 169–185. https://doi.org/10.1007/978-3-030-60939-9_12